OpenStudy (anonymous):

@Mathematics Using the discriminant, what values of c will make the quadratic y = 2x2 - 3x + c have one real solution? A)c = -9/8 B)c = 9/8 C)(9/8, ∞) D)[9/8, ∞)

OpenStudy (anonymous):

b^2 - 4ac=0 (-3)^2-4(2)c=0 9-8c=0 c=9/8

OpenStudy (anonymous):

discriminant for parabola 1) two real soloution if it is positive number 2) repeated root(one soloution) if it is 0 3) two complex root if it is negative